PLEASE HELP! Am on a maths revision website and just CANNOT work it out! i know it must be real simple, but can't see it! "What is q when: tq-r=3q+p" any help will be appriciated. Thankyou, Elliott. p.s. thankyou to "Stan" for helping me out yesterday!
Answer by stanbon(75887)

"What is q when: tq-r=3q+p"
Get the q terms on one side as follows:
tq-3q=r+p
Factor out the "q" to get:
q(t-3)=r+p
Solve for "q":
q=(r+p)/(t-3)
Cheers,
Stan H.